# Inferring Stellar Densities with Flexible Models I: The Distribution of RR Lyrae in the Milky Way with Gaia DR3

### v2

@@ 3 Hierarchical Bayesian Model of Multi-Gaussian Galaxy
The aim of this work is to find the maximum likelihood estimate for a given RR Lyrae density model described by parameters, $\boldsymbol{\theta}$, given our RR Lyrae distance catalog and selection function ($S(l,b,d)$). Specifically, we use a hierarchical Bayesian model to infer the true underlying stellar density distribution $n(\boldsymbol{x}\,;\,\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ are the parameters of our density model. We transform true 3D Galactocentric positions $\boldsymbol{x}$ to observed distances ($d$) and sky positions ($l,b$) and marginalize over the distance uncertainty as follows:
@@
To perform the transformation to and from Galactocentric positions, we assume a right-handed coordinate system with the Sun’s position at (8.25-8.25 kpc, 0, 20.8 pc; GRAVITY Collaboration et al., 2021; Leung et al., 2023). The measured sky positions ($l,b$) of are significantly more precise than the distances ($d$) so the integral over 3D position simply becomes a 1D integral over the distance. For computational efficiency, this integral is calculated via a Riemann sum from $x_{min}$ to $x_{max}$, which are defined as follows:
$d_{min}$ $=\min\begin{cases}d-3\sigma_{d}\\ 10^{-3}\end{cases}$ (3)
$d_{max}$ $=\max\begin{cases}d+3\sigma_{d}\\ 10^{6}\end{cases}$ (4)
@@
The $P(\boldsymbol{x}_{i}|\boldsymbol{\theta})$ term in Equation 1 is our density model of the Galaxy. In this work, we use a Gaussian Mixture Model (GMM) to describe the Galactic stellar density because it is a normalized probability density distribution unlike many of the power law models typically used to describe stellar halos (e.g., Deason et al., 2011; Iorio et al., 2018; Han et al., 2022). Furthermore, GMMs can accurately approximate many power law functions with only a few components (Cappellari, 2002; Hogg & Lang, 2013; Miller & van Dokkum, 2021).
We fit two GMMs which are then combined with weights that sum to 1 as our complete density model. All Gaussian components of our model are centered on (0,0,0) kpc corresponding to the Galactic center. One of the GMMs is set to model the inner Galaxy, while the other models the halo. For both, we set the covariance matrices so that they increase by a factor of two between components. The difference between the inner Galaxy and halo model lies in the scaling factor ($k$) applied to the covariance matrices. For the inner Galaxy GMM $k=0.1$ kpc, while $k=1.5$ kpc for the halo GMM. The inner Galaxy GMM has 5 components while the halo GMM has 6. They are set to overlap in that the smallest halo GMM component has $\sigma_{xy}$=1.5 kpc and the largest inner Galaxy component has $\sigma_{xy}=$1.6 kpc. Each GMM also includes its own flattening parameter ($q$) applied to the z-coordinate, which is nominally perpendicular to the Galactic disk. Therefore, for both GMMs, the non-rotated covariance matrix for the jthjth component is as follows:
@@
However, there is increasing evidence the the Galactic halo of the Milky Way may be rotated with respect to the Galactic disk (Iorio & Belokurov, 2019; Han et al., 2022; Nibauer & Bonaca, 2025). In other words, the Galactic halo may not be flattened in a direction directly perpendicular to the Galactic disk ($z$) but instead along another direction. To investigate this, we develop a rotated halo model. Specifically, we transform the halo components’ covariance matrices with the following rotation matrix:
@@
$\begin{split}n(\boldsymbol{x}_{i}|\boldsymbol{\theta})&=\\ &(1-A)\sum_{j=1}^{5}A_{H,j}\,\mathcal{N}(\boldsymbol{x}_{i}|\boldsymbol{0},R_{\rm rot}cov_{j}(k=2.0~\rm{kpc})R_{\rm rot}^{T})\\ &+A\sum_{j=1}^{5}A_{IG,j}\,\mathcal{N}(\boldsymbol{x}_{i}|\boldsymbol{0},cov_{j}(k=0.1~\rm{kpc}))\end{split}$ (7)
@@
We utilize the stick breaking transform to fit the amplitudes of the GMM components. For complete details of the transform we refer the reader to Stan Development Team (2024). In short, it transforms $J$ GMM amplitudes to J1J-1 unconstrained parameters. This leads to better convergence because it reduces the dimensionality and ensures the amplitudes sum to 1.
3.1 Inference
As our hierarchical Bayesian model is computationally expensive with high dimensionality, we require fast, scalable optimizers and posterior estimators. For fast maximum likelihood estimates, we use an the adafactor optimizer to find the minimum of the negative log-likelihood . We choose adafactor because it has similar benefits to the Adam optimizer (efficient stochastic gradient descent) with lower memory cost (Shazeer & Stern, 2018) which is crucial in this work given that all of the data has to be held in memory at once in order to calculate the full likelihood. To estimate the uncertainties, we calculate the covariance matrix at the maximum likelihood estimate (MLE) using the negative inverse Hessian.
4 The Galactic density distribution of RR Lyrae
@@
Table 1: Inferred Shapes and Tilt
Component q $\rm{\alpha}~(^{\circ})$ $\rm{\phi}~(^{\circ})$
Halo $0.697\pm 0.002$ $17.7\pm 0.7$ $27.2\pm 2.4$
Inner Galaxy $1.311\pm 0.002$ N/A N/A
@@
4.1 Comparison to Literature Halo Models
Figure 4: The normalized halo density profile of RR Lyrae as a function of distance from the Galactic center ($r$) for our GMM. The light purple dotted lines indicate the 5 components of the inner Galaxy GMM while the dark purple dotted lines show the halo GMM components. The dark blue solid line is their sum. For comparison, we also show results from the literature for halo profiles including Horta & Schiavon (2025, yellow), Han et al. (2022, green), Yang et al. (2022, light blue) and Pietrukowicz et al. (2015, red). The works from Horta & Schiavon (2025), Han et al. (2022), and Yang et al. (2022) use red giant stars while Pietrukowicz et al. (2015) utilize only RRL.
@@ 4.1 Comparison to Literature Halo Models
Galactic halos are commonly described with power-law profiles, including cored, (doubly) broken, triaxial, and rotated adaptations (Deason et al., 2011; Iorio et al., 2018; Han et al., 2022). However, in this work, we use a Gaussian Mixture model as an easily normalizable probability density and more flexible alternative to the power-law. This flexibility is especially crucial for modeling the inner Galaxy where the disk, halo and bulge all overlap.
We show the normalized density as as function of Galactocentric radius (r) for the halo components of our Gaussian Mixture Model Galaxy in Figure 4. Specifically, each of the 5 Gaussian halo components for the inner Galaxy GMM and the halo GMM are shown in light purple and dark purple dotted lines, respectively, with the sum shown in the solid dark blue line. Profiles from the literature are shown for comparison including results from Pietrukowicz et al. (2015, red), Yang et al. (2022, light blue), Han et al. (2022, green) and Horta & Schiavon (2025, yellow).
Each of the literature results shown use different stellar populations and methodologies to trace the stellar halo. For example, Pietrukowicz et al. (2015) uses RRL from the OGLE survey (Udalski et al., 2015) and converts the directly measured surface density to 3D density profile. Yang et al. (2022) estimates the halo profile using direct orbit integration of K giant stars with galactocentric distances between 5 and 50 kpc. They remove disk stars using a selection in [Fe/H] and cylindrical velocity. Similar to our method, Han et al. (2022) use a Bayesian model to infer the halo profile using a sample of stars selected to be from the Gaia-Enceladus-Sausage merger (Belokurov et al., 2018; Helmi et al., 2018) using $\alpha$/Fe] and [Fe/H] abundances as well as orbital eccentricity. Horta & Schiavon (2025) uses a similar method to Han et al. (2022), however they utilize a selection in [Mn/Mg], [Al/Fe], [Mg/Fe], and [Fe/H] to isolate the “Heracles/proto-galaxy" structure that is generally confined to the inner Galaxy with apocenters<12 kpc. Our work improves upon these previous results by not performing a chemical or dynamical selection and instead directly modeling all of the available data including stars with Galactocentric radii as small as $\approx$0.2 kpc.
@@
In addition to comparing the slope of our halo model to previous works, we can also compare the rotation. Specifically, we compare our model to the tilted halo model of Han et al. (2022), which was recently confirmed by an independent analysis in Nibauer & Bonaca (2025). In these works, they use a triaxial halo model, and relatedly, a different rotation matrix since their model does not have XY-symmetry, unlike ours. However, we can directly compare the pitch angles which both describe the tilt of the z-axis towards the positive x-axis. We find a pitch angle of $18^{\circ}$, in agreement with the halo mass distribution of Nibauer & Bonaca (2025) and stellar density distribution of Han et al. (2022). It is interesting to note that Han et al. (2022) specifically measure the stellar distribution of the Gaia-Sausage-Enceladus system, which makes up the bulk of the local halo (Belokurov et al., 2018; Helmi et al., 2018).
@@
To visually compare our rotated halo with that of Han et al. (2022), we plot density contours of the halo GMM in Figure 5, with the highest density in yellow and the lowest density in dark blue. The left panel shows the face-on slice evaluated at Z=0 kpc, while the middle and right panel show edge-on slices at X=0 kpc and Y=0 kpc, respectively. To guide the reader, we show the Sun’s position (black star) in the left and middle panels at X=-8.25 kpc. The shape and tilt of our halo model agrees well with results from Han et al. (2022) shown as a red dashed line, especially in the middle panel which shows the X-Z slice. They left and right panels show slight disagreement with the Han et al. (2022) results, but this is likely because of the difference in rotation matrices, where our model preserves X-Y symmetry while Han et al. (2022) does not.
@@
In Figure 5, we also show an arrow that points to the location of the Sagittarius dwarf galaxy at (17.9, 2.6, -6.6) kpc (Vasiliev et al., 2021). This algins well with the rotated major-axis of the halo models. Sagittarius is a clear non-axisymmetric feature in our observed data (see Figure 2) so it is not unexpected that it could cause the tilt in our inferred halo model. Han et al. (2022) targets modeling the halo contribution of only the Gaia-Enceladus-Sausage. Whether the tilt is caused by contamination from Sagittarius stars, or is a strange coincidence requires further investigation. Nibauer & Bonaca (2025) models the dark matter halo tilt, however they assume an axisymmetric subtraction of stellar mass that does not include Sagittarrius. This could cause the elongation of the “missing" mass model towards the Sagittarius system.
@@
For the inner Galaxy component, we measure an oblate shape with $q=1.31$. This is consistent with works measuring the shape of the dark matter halo within 20 kpc of the Galactic center using kinematic tracers (Bowden et al., 2016; Bovy et al., 2016a), including the work of Posti & Helmi (2019) which also found $q=1.3$. However, these works assume an oblate stellar halo or bulge for the non-disk component of the inner Galaxy. Further work to simultaneously study the stellar and dark matter mass distribution in the inner Galaxy is required to determine whether the RRL are tracing the distribution of dark matter as predicted by cosmological zoom-in simulations (Lucey et al., 2025).
5 Summary and Conclusions
@@
To address these observational challenges, we use the standard candle nature of RR Lyrae stars to map the structure of the Milky Way with a hierarchical Bayesian Gaussian Mixture Model (GMM). This is the first time a GMM has been used to describe the stellar density distribution of the Milky Way, building on the success of this method in characterizing external galaxies including the power laws typically used to describe halos (Cappellari, 2002; Hogg & Lang, 2013; Miller & van Dokkum, 2021). Crucially, we find the GMM provides the increased flexibility needed to model the inner Galaxy where the bulge, halo and disk all intersect.
We use a total of 10 Gaussian components to model the Galactic distribution of RRL. All of the components are centered on the Galactic center. Five of the components are set to model the inner Galaxy with XY-covariances between 0.1-1.6 kpc and allowing the covariance in the z-direction to differ by a factor of $q$. We use a similar setup for the 6 components that model the halo, but with larger covariances (1.5-48 kpc) and allowing the z-axis to be rotated with respect to the Galactic disk. This framework yields two key findings about the Galactic distribution of RR Lyrae:
  • The inner Galaxy is dominated by distinct prolate population with $q=1.31$, consistent with results for the shape of the dark matter distribution in the inner Galaxy (Bowden et al., 2016; Posti & Helmi, 2019).
@@
  • We infer a flattening (qq=0.70) and halo tilt in the RRL population ($\alpha=18^{\circ}$) consistent with Han et al. (2022) and Nibauer & Bonaca (2025). We note for the first time that the rotated major axis points to the Sagittarius dwarf galaxy.
These results demonstrate that GMMs offer a powerful and flexible framework for modeling complex stellar distributions in the Milky Way. The main limit to our work is the sample size, especially close to the Galactic plane where we are severely limited by dust extinction. This especially impacts the selection function resolution and precision. Larger samples of RRL with well-calibrated distances, especially in high extinction regions will greatly improve this work, which the future Roman Galactic Plane Survey should provide (Sanderson et al., 2024). Looking ahead, the next work in this series will apply this method to study how the halo structure varies as a function of stellar metallicity. Such studies will provide unprecedented insight into the formation history and chemical evolution of our Galaxy.

### v3

@@ III Hierarchical Bayesian Model of Multi-Gaussian Galaxy
The aim of this work is to find the maximum likelihood estimate for a given RRL density model described by parameters, $\boldsymbol{\theta}$, given our RRL distance catalog and selection function ($S(l,b,d)$). Specifically, we use a hierarchical Bayesian model to infer the true underlying stellar density distribution $n(\boldsymbol{x}\,;\,\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ are the parameters of our density model. We transform true 3D Galactocentric positions $\boldsymbol{x}$ to observed distances ($d$) and sky positions ($l,b$) and marginalize over the distance uncertainty as follows:
@@
To perform the transformation to and from Galactocentric positions, we assume a right-handed coordinate system with the Sun’s position at (8.25-8.25 kpc, 0, 20.8 pc; GRAVITY Collaboration et al., 2021; Leung et al., 2023). The measured sky positions ($l,b$) of are significantly more precise than the distances ($d$) so the integral over 3D position simply becomes a 1D integral over the distance. For computational efficiency, this integral is calculated via a Riemann sum from $d_{min}$ to $d_{max}$, which are defined as follows:
$d_{min}$ $=\min\begin{cases}d-3\sigma_{d}~[\rm{kpc}]\\ 10^{-3}~\rm{kpc}\end{cases}$ (3)
$d_{max}$ $=d+3\sigma_{d}~[\rm{kpc}]$ (4)
@@
The $P(\boldsymbol{x}_{i}|\boldsymbol{\theta})$ term in Equation 1 is our density model of the Galaxy. In this work, we use a GMM to describe the Galactic stellar density because it is a normalized probability density distribution unlike many of the power law models typically used to describe stellar halos (e.g., Deason et al., 2011; Iorio et al., 2018; Han et al., 2022). Furthermore, GMMs can accurately approximate many power law functions with only a few components (Cappellari, 2002; Hogg and Lang, 2013; Miller and van Dokkum, 2021). While the choice of GMM is not physically-motivated, it offers numerous mathematical advantages (Bendinelli, 1991; Bovy et al., 2011), including an analytical solution to the Jean’s equations (Cappellari, 2008), and better fits to real data where simple parametric models fail to recover the full complexity of galactic structure (Emsellem et al., 1994).
We fit two GMMs which are then combined with weights that sum to 1 as our complete density model. All Gaussian components of our model are centered on (0,0,0) kpc corresponding to the Galactic center. One of the GMMs is set to model the inner Galaxy, while the other models the halo. For both, we set the covariance matrices so that they increase by a factor of two between components. The difference between the inner Galaxy and halo model lies in the scaling factor ($k$) applied to the covariance matrices. For the inner Galaxy GMM $k=0.1$ kpc, while $k=1.5$ kpc for the halo GMM. The inner Galaxy GMM has 5 components while the halo GMM has 6. Using 6–11 total GMM components with logarithmically scaled widths is standard for modeling galaxy profiles (Cappellari, 2002; Hogg and Lang, 2013; Miller and van Dokkum, 2021). Our method adds additional flexibility in allowing the halo to be flattened in a direction different than in the inner Galaxy. To allow flexibility in the model for exactly where the halo GMM begins to dominate the density profile, we ensure overlap in that the smallest halo GMM component has $\sigma_{xy}$=1.5 kpc and the largest inner Galaxy component has $\sigma_{xy}=$1.6 kpc. Each GMM also includes its own flattening parameter ($q$) applied to the z-coordinate, which is nominally perpendicular to the Galactic disk. Therefore, for both GMMs, the non-rotated covariance matrix for the jthjth component is as follows:
@@
However, there is increasing evidence the the Galactic halo of the Milky Way may be rotated with respect to the Galactic disk (Iorio and Belokurov, 2019; Han et al., 2022; Nibauer and Bonaca, 2025). In other words, the Galactic halo may not be flattened in a direction directly perpendicular to the Galactic disk ($z$) but instead along another direction. To investigate this, we develop a rotated halo model. Specifically, we transform the halo components’ covariance matrices with the following rotation matrix:
@@
$\begin{split}n(\boldsymbol{x}_{i}|\boldsymbol{\theta})&=\\ &(1-A)\sum_{j=1}^{6}A_{H,j}\,\mathcal{N}(\boldsymbol{x}_{i}|\boldsymbol{0},R_{\rm rot}cov_{j}(k=2.0~\rm{kpc})R_{\rm rot}^{T})\\ &+A\sum_{j=1}^{5}A_{IG,j}\,\mathcal{N}(\boldsymbol{x}_{i}|\boldsymbol{0},cov_{j}(k=0.1~\rm{kpc}))\end{split}$ (7)
@@
We utilize the stick breaking transform to fit the amplitudes of the GMM components. For complete details of the transform we refer the reader to Stan Development Team (2024). In short, it transforms $J$ GMM amplitudes to J1J-1 unconstrained parameters. This leads to better convergence because it reduces the dimensionality and ensures the amplitudes sum to 1. With this transform, our model has a total of 14 free parameters (10 amplitudes, 2 flattening parameters ($q$), and 2 angles) to fit.
III.1 Inference
As our hierarchical Bayesian model is computationally expensive with high dimensionality (14 free parameters), we require fast, scalable optimizers and posterior estimators. For fast maximum likelihood estimates, we use an the adafactor optimizer to find the minimum of the negative log-likelihood . We choose adafactor because it has similar benefits to the Adam optimizer (efficient stochastic gradient descent) with lower memory cost (Shazeer and Stern, 2018) which is crucial in this work given that all of the data has to be held in memory at once in order to calculate the full likelihood. To estimate the uncertainties, we calculate the covariance matrix at the maximum likelihood estimate (MLE) using the negative inverse Hessian.
IV The Galactic density distribution of RR Lyrae
@@
Table 1: Inferred Parameters of GMM Model
Inner Galaxy Halo
q $1.311\pm 0.002$ $0.697\pm 0.002$
$\rm{\alpha}~(^{\circ})$ N/A $17.7\pm 0.7$
$\rm{\phi}~(^{\circ})$ N/A $27.2\pm 2.4$
Amplitudes for GMM
A 0.37
$j=$ $A_{IG,j}$ $A_{H,j}$
1 $7.00\times 10^{-2}$ $9.99\times 10^{-5}$
2 $6.53\times 10^{-2}$ $1.14\times 10^{-4}$
3 $6.03\times 10^{-2}$ $5.60\times 10^{-2}$
4 $5.05\times 10^{-2}$ $7.70\times 10^{-1}$
5 $7.53\times 10^{-1}$ $1.32\times 10^{-1}$
6 $4.25\times 10^{-2}$
@@
IV.1 Comparison to Literature Halo Models
Figure 4: The normalized halo density profile of RRL as a function of distance from the Galactic center ($r$) for our GMM compared to literature results using RRL (left) and other tracers (right). In both panels, the light purple dotted lines indicate the 5 components of the inner Galaxy mixture model while the dark purple dotted lines show the halo components. The dark blue solid line is their sum. The literature RRL profiles included in the left panel are Pietrukowicz et al. (2015, red), Iorio et al. (2018, blue) and Medina et al. (2024, green). The works in the right panel, Horta and Schiavon (2025, yellow), Han et al. (2022, green), and Yang et al. (2022, light blue), use red giant stars.
@@ IV.1 Comparison to Literature Halo Models
Galactic halos are commonly described with power-law profiles, including cored, (doubly) broken, triaxial, and rotated adaptations (Deason et al., 2011; Iorio et al., 2018; Han et al., 2022). However, in this work, we use a GMM as an easily normalizable probability density and more flexible alternative to the power-law. This flexibility is especially crucial for modeling the inner Galaxy where the disk, halo and bulge all overlap.
We show the normalized density as as function of Galactocentric radius (r) for the halo components of our GMM Galaxy in Figure 4, compared to results from the literature using RRL in the left panel and other tracers in the right panel. In each panel, the 5 Gaussian halo components for the inner Galaxy GMM and the halo GMM are shown in light purple and dark purple dotted lines, respectively, with the sum shown in the solid dark blue line. In the left panel, the RRL profiles from the literature include results from Pietrukowicz et al. (2015, red), Iorio et al. (2018, blue), and Medina et al. (2024, green). These results are chosen to cover a wide range of radii, and includes both single and double power-law results, prioritizing recent work. The right panels includes the works of Yang et al. (2022, light blue), Han et al. (2022, green) and Horta and Schiavon (2025, yellow).
The literature results shown in the right panel are chosen to highlight the diversity of profiles measured using different stellar populations and methodologies to trace the stellar halo. For example, Yang et al. (2022) estimates the halo profile using direct orbit integration of K giant stars with galactocentric distances between 5 and 50 kpc. They remove disk stars using a selection in [Fe/H] and cylindrical velocity. Similar to our method, Han et al. (2022) use a Bayesian model to infer the halo profile using a sample of stars selected to be from the Gaia-Enceladus-Sausage merger (Belokurov et al., 2018; Helmi et al., 2018) using $\alpha$/Fe] and [Fe/H] abundances as well as orbital eccentricity. Horta and Schiavon (2025) uses a similar method to Han et al. (2022), however they utilize a selection in [Mn/Mg], [Al/Fe], [Mg/Fe], and [Fe/H] to isolate the “Heracles/proto-galaxy" structure that is generally confined to the inner Galaxy with apocenters<12 kpc. Our work improves upon these previous results by not performing a chemical or dynamical selection and instead directly modeling all of the available data including stars with Galactocentric radii as small as $\approx$0.2 kpc.
@@
In addition to comparing the slope of our halo model to previous works, we can also compare the rotation. Specifically, we compare our model to the tilted halo model of Han et al. (2022), which was recently confirmed by an independent analysis in Nibauer and Bonaca (2025). In these works, they use a triaxial halo model, and relatedly, a different rotation matrix since their model does not have XY-symmetry, unlike ours. However, we can directly compare the pitch angles which both describe the tilt of the z-axis towards the positive x-axis. We find a pitch angle of $18^{\circ}$, in agreement with the halo mass distribution of Nibauer and Bonaca (2025) and stellar density distribution of Han et al. (2022). It is interesting to note that Han et al. (2022) specifically measure the stellar distribution of the Gaia-Sausage-Enceladus system, which makes up the bulk of the local halo (Belokurov et al., 2018; Helmi et al., 2018).
@@
To visually compare our rotated halo with that of Han et al. (2022), we plot density contours of the halo GMM in Figure 5, with the highest density in yellow and the lowest density in dark blue. The left panel shows the face-on slice evaluated at Z=0 kpc, while the middle and right panel show edge-on slices at X=0 kpc and Y=0 kpc, respectively. To guide the reader, we show the Sun’s position (black star) in the left and middle panels at X=-8.25 kpc. The shape and tilt of our halo model agrees well with results from Han et al. (2022) shown as a red dashed line, especially in the middle panel which shows the X-Z slice. The left and right panels show slight disagreement with the Han et al. (2022) results, but this is likely because of the difference in rotation matrices, where our model preserves X-Y symmetry while Han et al. (2022) does not.
@@
In Figure 5, we also show an arrow that points to the location of the Sagittarius dwarf galaxy at (17.9, 2.6, -6.6) kpc (Vasiliev et al., 2021). This aligns well with the rotated major-axis of the halo models. Sagittarius is a clear non-axisymmetric feature in our observed data (see Figure II) so it is not unexpected that it could cause the tilt in our inferred halo model. Han et al. (2022) targets modeling the halo contribution of only the Gaia-Enceladus-Sausage. Whether the tilt is caused by contamination from Sagittarius stars, or is a strange coincidence requires further investigation, including ongoing theoretical work testing the impact of assuming a single halo model when a galaxy is undergoing a Sagittarius-like merger versus explicitly modeling the second, less massive galaxy. Nibauer and Bonaca (2025) models the dark matter halo tilt, however they assume an axisymmetric subtraction of stellar mass that does not include Sagittarius. This could cause the elongation of the “missing" mass model towards the Sagittarius system.
@@
For the inner Galaxy component, we measure an oblate shape with $q=1.31$. This is consistent with works measuring the shape of the dark matter halo within 20 kpc of the Galactic center using kinematic tracers (Bowden et al., 2016; Bovy et al., 2016a), including the work of Posti and Helmi (2019) which also found $q=1.3$. However, these works assume an oblate stellar halo or bulge for the non-disk component of the inner Galaxy. Further work to simultaneously study the stellar and dark matter mass distribution in the inner Galaxy is required to determine whether the RRL are tracing the distribution of dark matter as predicted by cosmological zoom-in simulations (Lucey et al., 2025).
V Summary and Conclusions
@@
To address these observational challenges, we use the standard candle nature of RRL stars to map the structure of the Milky Way with a hierarchical Bayesian GMM. This is the first time a GMM has been used to describe the stellar density distribution of the Milky Way, building on the success of this method in characterizing external galaxies including the power laws typically used to describe halos (Cappellari, 2002; Hogg and Lang, 2013; Miller and van Dokkum, 2021). Crucially, we find the GMM provides the increased flexibility needed to model the inner Galaxy where the bulge, halo and disk all intersect.
We use a total of 11 Gaussian components to model the Galactic distribution of RRL. All of the components are centered on the Galactic center. Five of the components are set to model the inner Galaxy with XY-covariances between 0.1-1.6 kpc and allowing the covariance in the z-direction to differ by a factor of $q$. We use a similar setup for the 6 components that model the halo, but with larger covariances (1.5-48 kpc) and allowing the z-axis to be rotated with respect to the Galactic disk. This framework yields three key findings about the Galactic distribution of RRL:
  • The inner Galaxy $(R\lesssim 10~\rm{kpc})$ is dominated by distinct prolate population with $q=1.31$, consistent with results for the shape of the dark matter distribution in the inner Galaxy (Bowden et al., 2016; Posti and Helmi, 2019).
@@
  • We infer a flattening (qq=0.70) and halo tilt in the RRL population ($\alpha=18^{\circ}$) consistent with Han et al. (2022) and Nibauer and Bonaca (2025). We note for the first time that the rotated major axis points to the Sagittarius dwarf galaxy.
These results demonstrate that GMMs offer a powerful and flexible framework for modeling complex stellar distributions in the Milky Way. The main limit to our work is the sample size, especially close to the Galactic plane where we are severely limited by dust extinction. This especially impacts the selection function resolution and precision. Larger samples of RRL with well-calibrated distances, especially in high extinction regions will greatly improve this work, which the future Roman Galactic Plane Survey should provide (Sanderson et al., 2024). Looking ahead, the next work in this series will apply this method to study how the halo structure varies as a function of stellar metallicity. As the relationship between the stellar halo’s metallicity and structure is indicative of its origins (Eggen et al., 1962; Searle and Zinn, 1978), such studies will provide insight into the formation history and chemical evolution of our Galaxy.